# Judging with Many Minds: Do More Perspectives Mean Less Prejudice? On Bias Amplifications and Resistance in Multi-Agent Based LLM-as-Judge

### v2

@@ Abstract
LLM-as-Judge has emerged as a scalable alternative to human evaluation, enabling large language models (LLMs) to provide reward signals in trainings. While recent work has explored multi-agent extensions such as multi-agent debate and meta-judging to enhance evaluation quality, the question of how intrinsic biases manifest in these settings remains underexplored. In this study, we conduct a systematic analysis of four diverse bias types: position bias, verbosity bias, chain-of-thought bias, and bandwagon bias. We evaluate these biases across two widely adopted multi-agent LLM-as-Judge frameworks: Multi-Agent-Debate and LLM-as-Meta-Judge. Our results show that debate framework amplifies biases sharply after the initial debate, and this increased bias is sustained in subsequent rounds, while meta-judge approaches exhibit greater resistance. We further investigate the incorporation of PINE, a leading single-agent debiasing method, as a bias-free agent within these systems. The results reveal that this bias-free agent effectively reduces biases in debate settings but provides less benefit in meta-judge scenarios. Our work provides a comprehensive study of bias behavior in multi-agent LLM-as-Judge systems and highlights the need for targeted bias mitigation strategies in collaborative evaluation settings.
Judging with Many Minds: Do More Perspectives Mean Less Prejudice? On Bias Amplifications and Resistance in Multi-Agent Based LLM-as-Judge
@@ 3.2 Multi-Agent-Debate
In the Multi-Agent-Debate setting, we adopt and implement the “General Public and Critic” structure, as originally proposed in ChatEval Chan et al. (2023), and widely used in several recent studies Feng et al. (2024); Arif et al. (2024); Zhang et al. (2025a). The General Public agent is analogous to the judge model in the single-agent LLM-as-Judge setting. It is tasked with making an initial judgment, which includes the scores and reasoning for both solutions. The Critic agent, in contrast, first reads and comprehends the evaluation task and the judgment provided by the General Public agent. It then offers its critique of the initial judgment along with its own independent assessment. To mirror real-world debate scenarios, the Critic agent is encouraged to provide a different judgment from that of the General Public agent. In subsequent rounds of conversation, the General Public agent must consider the Critic’s feedback and may choose either to revise or to maintain its original judgment. The Critic agent continues to generate additional critiques and independent assessments based on the evolving conversation. At the end of the debate, we use the judgment of the General Public agent as the final verdict. A demonstration of the debate flow is shown in the left panel of Figure 1. The prompts used for the General Public and Critic agents are provided in Appendix D.1.
@@ 3.3 LLM-as-Meta-Judge
In the LLM-as-Meta-Judge setting, we focus on the fundamental concept of “Thinking about thinking.” Inspired by the frameworks proposed by Li et al. and Wu et al., we adopt diverse LLMs as judges, each providing an independent judgment to form a judgment pool based on the evaluation task. We further examine two variations of the Meta-Judge framework, as shown in the right panel of Figure 1. In the “Select” setting, the Meta-Judge selects the best judgment from the pool and uses it as the final verdict. This setting evaluates the Meta-Judge’s ability to discern and identify the highest-quality answer among diverse alternatives, testing its competence as an evaluator and ranker of peer outputs. In contrast, the “Conclude” setting requires the Meta-Judge to provide its own judgment as the final verdict based on the judgment pool. This setting assesses the Meta-Judge’s ability to integrate, reason over, and reconcile potentially conflicting or complementary information, acting as a collaborative decision-maker. It is also similar to the “summarization” role that was widely used in group discussion settings Liu et al. (2024c); Li et al. (2025b). The prompts used for the Meta Judges are provided in Appendix D.2.
@@ 3.4 Biases and Measurement
In our study, we primarily focus on four types of intrinsic bias: position bias, verbosity bias, bandwagon bias, and chain-of-thought bias. Illustrative examples of these biases are provided in Figure 2. We selected these biases because they span some of the most critical and generalizable dimensions of intrinsic model behavior: input ordering, response length, reasoning stability, and social conformity. To measure them, we follow the procedures proposed by CALM Ye et al. (2024), which assess bias by comparing the consistency of judgments before and after specific prompt modifications designed to elicit each bias. In the Multi-Agent LLM-as-Judge framework, all agents are exposed to these prompt modifications when measuring biases. Examples of these prompts can be found in Appendix D.3. The consistency rate is defined as:
@@ 4 Experiment
$Pos.$ $Verbo.$ $Band.$ $Cot$ $Pos.$ $Verbo.$ $Band.$ $Cot$
@@ 4.1 Evaluation Dataset
We conduct our experiments on two benchmarks: MT-Bench Zheng et al. (2023) and the Alignment dataset from CALM Ye et al. (2024), which we refer to as CALM-Alignment. MT-Bench consists of responses from 30 LLMs to each of 8 tasks, with 10 questions per task. For each question, we sample 60 response pairs per task along with their corresponding questions, yielding a total of 480 samples. This sampling strategy balances computational cost with comparability to the CALM-Alignment dataset. CALM-Alignment, sampled from various DPO datasets, contains 439 samples. Each response pair includes a chosen and a rejected label based on human feedback. More details on the sub-sampled MT-Bench and CALM-Alignment datasets are provided in Appendix B. We apply prompt modifications as described in Sec. 3.4. Notably, since the CALM-Alignment dataset includes full labels, we specifically extend the solution for the rejected responses (which always appear as the second solution in the pair in original order). This allows us to measure how many previously correct judgments are changed to incorrect ones. For consistency, we also apply the prompt extension to the second solution in each pair in MT-Bench.
@@ 4.2 Experiments Setup
To assess the impact of a bias-free agent, we apply the PINE method Wang et al. (2024b) to Qwen1.5-32B-Chat Bai et al. (2023), strictly following the setup used in the original work. We then deploy it in debates against GPT-4o-mini (which exhibited the largest drop in position bias consistency) and LLaMA 3.3 70B (which showed the least drop). Leveraging PINE’s position bias elimination, we investigate whether a bias-free agent can reduce bias in judge models through interaction. In our experiments, the bias-free agent serves as a critic in the Multi-Agent Debate and is included in the judgment pool under Meta-Judge settings. This experiment targets scenarios where modifying the judge model or its prompt is impractical, such as self-reward settings during training Zhang et al. (2024); Wu et al. (2024). Full results are presented in Table 3 and Table 4. Case studies and examples can be found in Appendix E.
@@
R1-Distilled-Qwen 32B 4o-mini Llama3.3 70B DeepSeek-V3 $Mode$ $Pos.$ $Verb.$ $Band.$ $Cot.$ $Pos.$ $Verb.$ $Band.$ $Cot.$
@@
R1-Distilled-Qwen 32B 4o-mini Llama3.3 70B DeepSeek-V3 $Mode$ $Pos.$ $Verb.$ $Band.$ $Cot.$ $Pos.$ $Verb.$ $Band.$ $Cot.$
@@
Meta Judge: DeepSeek Chat
R1-Distilled-Qwen 32B 4o-mini Llama3.3 70B DeepSeek-V3 $Mode$ $Pos.$ $Verb.$ $Band.$ $Cot.$ $Pos.$ $Verb.$ $Band.$ $Cot.$
@@ 5 Results and Analysis
Meta Judge: DeepSeek Chat
@@
Meta Judge: Llama3.3 70B
@@ 5.3 Bias-Free Experiment
Although the bias-free agent performs well in debate settings, it has minimal effect when used as a Meta-Judge. As shown in the Table 3, we observe a continuous improvement in position bias consistency with increasing rounds of debate. This is significantly different from the what we observed without such bias-free agent. On the other hand, as shown in Table 4, there is no considerable improvement to the consistency score when including this bias-free agent into the judgment pool. Our analysis further shows that, in select mode, the Meta-Judge shows no favor to the bias-free agent. This might suggest that the Meta-Judge itself may not be sensitive enough to recognize or value the unbiased nature of the agent’s outputs, instead treating all pool members as equally valid. As a result, the bias-free agent’s potential for debiasing is diluted within the collective judgment process. More details and analysis can be found in Appendix A.3.
@@ A.3 Analysis of Meta-Judge Selections with PINE
Appendix B More information on Benchmarks
B.1 MT-Bench
@@
The first benchmark that we use in our work is MT-Bench Zheng et al. (2023), which originally contains responses from 30 LLMs to 80 questions. The questions are divided into eight categories: Writing, Roleplay, Reasoning, Math, Coding, Extraction, STEM, and Humanities, with ten questions in each category. The LLMs include GPT-4 Turbo, Claude Instant v1, Vicuna 13B v1.3, Alpaca 13B, and others. To construct our evaluation set, we treat each question paired with two distinct model responses as a single pairwise comparison task. For each question, we generate all possible unique response pairs between the selected LLMs, ensuring no repetition. From this set, we randomly sample 60 response pairs per category, each coupled with its corresponding question. This procedure yields a total of 480 comparison tasks across the eight categories.
MT-bench is public available with License Apache 2.0.
B.2 CALM Benchmark
@@
Appendix C More information on PINE
@@
Appendix D Prompt Examples
D.1 Multi-Agent-Debate
@@
D.2 LLM-as-Meta Judge Prompts
@@ D.3 Bias Prompts
Appendix E Case Studies of Multi-Agent Debate
E.1 Original Debate
@@ E.3 With PINE
Appendix F Case Studies of LLM-as-Meta-Judge
F.1 Original Debate
@@ F.3 With PINE
Appendix G Implementation and Reproducibility
@@
To ensure the reproducibility of our experiments, we set fixed random seeds in our codes. All language models are evaluated with a temperature of $\text{temperature}=0.01$ to minimize randomness in their responses. The codes are available at https://github.com/Henrymachiyu/Multi_Agent_Judge_Bias

### v3

# Judging with Many Minds: Do More Perspectives Mean Less Prejudice? On Bias Amplification and Resistance in Multi-Agent Based LLM-as-Judge

@@ Abstract
LLM-as-Judge has emerged as a scalable alternative to human evaluation, enabling large language models (LLMs) to provide reward signals in trainings. While recent work has explored multi-agent extensions such as multi-agent debate and meta-judging to enhance evaluation quality, the question of how intrinsic biases manifest in these settings remains underexplored. In this study, we conduct a systematic analysis of four diverse bias types: position bias, verbosity bias, chain-of-thought bias, and bandwagon bias. We evaluate these biases across two widely adopted multi-agent LLM-as-Judge frameworks: Multi-Agent-Debate and LLM-as-Meta-Judge. Our results show that debate framework amplifies biases sharply after the initial debate, and this increased bias is sustained in subsequent rounds, while meta-judge approaches exhibit greater resistance. We further investigate the incorporation of PINE, a leading single-agent debiasing method, as a bias-free agent within these systems. The results reveal that this bias-free agent effectively reduces biases in debate settings but provides less benefit in meta-judge scenarios. Our work provides a comprehensive study of bias behavior in multi-agent LLM-as-Judge systems and highlights the need for targeted bias mitigation strategies in collaborative evaluation settings. Our code and experimental data can be found at https://github.com/Henrymachiyu/Multi_Agent_Judge_Bias.
Judging with Many Minds: Do More Perspectives Mean Less Prejudice? On Bias Amplification and Resistance in Multi-Agent Based LLM-as-Judge
@@ 3.2 Multi-Agent-Debate
In the Multi-Agent-Debate setting, we adopt and implement the “General Public and Critic” structure, as originally proposed in ChatEval Chan et al. (2023), and widely used in several recent studies Feng et al. (2024); Arif et al. (2024); Zhang et al. (2025a). The General Public agent is analogous to the judge model in the single-agent LLM-as-Judge setting. It is tasked with making an initial judgment, which includes the scores and reasoning for both solutions. The Critic agent, in contrast, first reads and comprehends the evaluation task and the judgment provided by the General Public agent. It then offers its critique of the initial judgment along with its own independent assessment. To mirror real-world debate scenarios, the Critic agent is encouraged to provide a different judgment from that of the General Public agent. In subsequent rounds of conversation, the General Public agent must consider the Critic’s feedback and may choose either to revise or to maintain its original judgment. The Critic agent continues to generate additional critiques and independent assessments based on the evolving conversation. At the end of the debate, we use the judgment of the General Public agent as the final verdict. A demonstration of the debate flow is shown in the left panel of Figure 1. The prompts used for the General Public and Critic agents are provided in Appendix F.1.
@@ 3.3 LLM-as-Meta-Judge
In the LLM-as-Meta-Judge setting, we focus on the fundamental concept of “Thinking about thinking.” Inspired by the frameworks proposed by Li et al. and Wu et al., we adopt diverse LLMs as judges, each providing an independent judgment to form a judgment pool based on the evaluation task. We further examine two variations of the Meta-Judge framework, as shown in the right panel of Figure 1. In the “Select” setting, the Meta-Judge selects the best judgment from the pool and uses it as the final verdict. This setting evaluates the Meta-Judge’s ability to discern and identify the highest-quality answer among diverse alternatives, testing its competence as an evaluator and ranker of peer outputs. In contrast, the “Conclude” setting requires the Meta-Judge to provide its own judgment as the final verdict based on the judgment pool. This setting assesses the Meta-Judge’s ability to integrate, reason over, and reconcile potentially conflicting or complementary information, acting as a collaborative decision-maker. It is also similar to the “summarization” role that was widely used in group discussion settings Liu et al. (2024c); Li et al. (2025b). The prompts used for the Meta Judges are provided in Appendix F.2.
@@ 3.4 Biases and Measurement
In our study, we primarily focus on four types of intrinsic bias: position bias, verbosity bias, bandwagon bias, and chain-of-thought bias. Illustrative examples of these biases are provided in Figure 2. We selected these biases because they span some of the most critical and generalizable dimensions of intrinsic model behavior: input ordering, response length, reasoning stability, and social conformity. To measure them, we follow the procedures proposed by CALM Ye et al. (2024), which assess bias by comparing the consistency of judgments before and after specific prompt modifications designed to elicit each bias. In the Multi-Agent LLM-as-Judge framework, all agents are exposed to these prompt modifications when measuring biases. Examples of these prompts can be found in Appendix F.3. The consistency rate is defined as:
@@ 4 Experiment
$Pos.$ $Verbo.$ $Band.$ CotCot $Pos.$ $Verbo.$ $Band.$ CotCot
@@ 4.1 Evaluation Dataset
We conduct our experiments on two benchmarks: MT-Bench Zheng et al. (2023) and the Alignment dataset from CALM Ye et al. (2024), which we refer to as CALM-Alignment. MT-Bench consists of responses from 30 LLMs to each of 8 tasks, with 10 questions per task. For each question, we sample 60 response pairs per task along with their corresponding questions, yielding a total of 480 samples. This sampling strategy balances computational cost with comparability to the CALM-Alignment dataset. CALM-Alignment, sampled from various DPO datasets, contains 439 samples. Each response pair includes a chosen and a rejected label based on human feedback. More details on the sub-sampled MT-Bench and CALM-Alignment datasets are provided in Appendix D. We apply prompt modifications as described in Sec. 3.4. Notably, since the CALM-Alignment dataset includes full labels, we specifically extend the solution for the rejected responses (which always appear as the second solution in the pair in original order). This allows us to measure how many previously correct judgments are changed to incorrect ones. For consistency, we also apply the prompt extension to the second solution in each pair in MT-Bench. To further support our findings, we also include results on Reward Bench Lambert et al. (2024) in Appendix B.
@@ 4.2 Experiments Setup
To assess the impact of a bias-free agent, we apply the PINE method Wang et al. (2024b) to Qwen1.5-32B-Chat Bai et al. (2023), strictly following the setup used in the original work. We then deploy it in debates against GPT-4o-mini (which exhibited the largest drop in position bias consistency) and LLaMA 3.3 70B (which showed the least drop). Leveraging PINE’s position bias elimination, we investigate whether a bias-free agent can reduce bias in judge models through interaction. In our experiments, the bias-free agent serves as a critic in the Multi-Agent Debate and is included in the judgment pool under Meta-Judge settings. This experiment targets scenarios where modifying the judge model or its prompt is impractical, such as self-reward settings during training Zhang et al. (2024); Wu et al. (2024). Full results are presented in Table 3 and Table 4. Case studies and examples can be found in Appendix G.
@@
R1-Distilled-Qwen 32B 4o-mini Llama3.3 70B DeepSeek-V3 ModeMode $Pos.$ $Verb.$ $Band.$ $Cot.$ $Pos.$ $Verb.$ $Band.$ $Cot.$
@@
R1-Distilled-Qwen 32B 4o-mini Llama3.3 70B DeepSeek-V3 ModeMode $Pos.$ $Verb.$ $Band.$ $Cot.$ $Pos.$ $Verb.$ $Band.$ $Cot.$
@@
Meta Judge: DeepSeek-V3
R1-Distilled-Qwen 32B 4o-mini Llama3.3 70B DeepSeek-V3 ModeMode $Pos.$ $Verb.$ $Band.$ $Cot.$ $Pos.$ $Verb.$ $Band.$ $Cot.$
@@ 5 Results and Analysis
Meta Judge: Llama3.3 70B
@@
Meta Judge: DeepSeek-V3
@@ 5.3 Bias-Free Experiment
Although the bias-free agent performs well in debate settings, it has minimal effect when used as a Meta-Judge. As shown in the Table 3, we observe a continuous improvement in position bias consistency with increasing rounds of debate. This is significantly different from the what we observed without such bias-free agent. On the other hand, as shown in Table 4, there is no considerable improvement to the consistency score when including this bias-free agent into the judgment pool. Our analysis further shows that, in select mode, the Meta-Judge shows no favor to the bias-free agent. This might suggest that the Meta-Judge itself may not be sensitive enough to recognize or value the unbiased nature of the agent’s outputs, instead treating all pool members as equally valid. As a result, the bias-free agent’s potential for debiasing is diluted within the collective judgment process. More details and analysis can be found in Appendix A.3. More results of PINE with Llama3 as base model can be found in Appendix C.
@@ 6 Conclusion
Acknowledgment
This research was supported in part by a generous grant from the John Templeton Foundation.
@@ A.3 Analysis of Meta-Judge Selections with PINE
Appendix B Results on RewardBench
To further support the observations we maded earlier, we random sample of 400 instances from the RewardBenchLambert et al. (2024) and perform our test on both Multi-Agent-Debate and LLM-as-Meta-Judge framework. We choose the sample size of 400 to specifically match the sample size that we used earlier for the experiments. The full results could be found in Table 14 for Multi-Agent-Debate, and Table 15 for LLM-as-Meta-Judge. The results on this benchmark are consistent with those from the initial two, further supporting the robustness of our findings across different settings. Specifically, biases tend to amplify sharply in the first round of debates and then stabilize in later rounds. On the other hand, expanding the judge pool shows more beneficial bias mitigation for Meta-Judge settings, consistent with our earlier observations.
Judge Critic Round Pos. Verbo. Band. Cot.
4o-mini 4o-mini 0 0.870 0.787 0.862 0.862
4o-mini 4o-mini 1 0.777 0.710 0.805 0.775
4o-mini 4o-mini 2 0.785 0.695 0.782 0.740
4o-mini 4o-mini 3 0.777 0.695 0.802 0.747
Llama 3.3 70B Llama 3.3 70B 0 0.871 0.834 0.872 0.909
Llama 3.3 70B Llama 3.3 70B 1 0.845 0.775 0.840 0.837
Llama 3.3 70B Llama 3.3 70B 2 0.850 0.785 0.850 0.832
Llama 3.3 70B Llama 3.3 70B 3 0.827 0.750 0.862 0.850
DeepSeek-V3 DeepSeek-V3 0 0.912 0.867 0.862 0.927
DeepSeek-V3 DeepSeek-V3 1 0.837 0.810 0.820 0.867
DeepSeek-V3 DeepSeek-V3 2 0.820 0.802 0.820 0.857
DeepSeek-V3 DeepSeek-V3 3 0.827 0.820 0.822 0.870
Table 14: Consistency scores for each bias (Positivity, Verbosity, Bandwagon, Cot) across three rounds of multi-agent debate, shown for different Judge–Critic model pairs on a subset of RewardBench.
Type Judge Referees Pos. Verbo. Band. Cot.
Judge 4o-mini - 0.870 0.787 0.862 0.862
Choose 4o-mini R1-Distilled-Qwen32B, 4o-mini 0.895 0.820 0.865 0.895
Choose 4o-mini R1-Distilled-Qwen32B, Llama3.3 70B 0.885 0.840 0.852 0.900
Choose 4o-mini R1-Distilled-Qwen32B, DeepSeek-V3 0.900 0.855 0.845 0.912
Choose 4o-mini R1-Distilled-Qwen32B, DeepSeek-V3, Llama3.3 70B 0.902 0.845 0.862 0.915
Choose 4o-mini R1-Distilled-Qwen32B, 4o-mini, DeepSeek-V3, Llama3.3 70B 0.895 0.835 0.875 0.902
Conclude 4o-mini R1-Distilled-Qwen32B, 4o-mini 0.867 0.805 0.857 0.867
Conclude 4o-mini R1-Distilled-Qwen32B, Llama3.3 70B 0.860 0.847 0.882 0.917
Conclude 4o-mini R1-Distilled-Qwen32B, DeepSeek-V3 0.867 0.827 0.847 0.900
Conclude 4o-mini R1-Distilled-Qwen32B, DeepSeek-V3, Llama3.3 70B 0.882 0.840 0.857 0.905
Conclude 4o-mini R1-Distilled-Qwen32B, 4o-mini, DeepSeek-V3, Llama3.3 70B 0.887 0.847 0.892 0.895
Table 15: Consistency scores for LLM-as-Meta-Judge across biases, pool sizes, and modes. "Yes" indicates inclusion of the model in the judgment pool; "No" indicates exclusion (single judge).
Appendix C More Results on Bias Free Agent
As discussed in the Limitations section, we implemented PINE on Qwen 1.5 to align with the original PINE studyWang et al. (2024b), which was designed for older models. While newer models may or may not offer better performance, our study is not focused on benchmarking the bias-free agent. Rather, we aim to show that ideal bias-free methods in single-agent settings do not fully mitigate biases in collaborative, multi-agent scenarios. However, to complement our study, we also tested Llama 3 8B-instruct (as proposed in the original PINE study) on the MT-Bench data. We found a consistent improvement in bias mitigation for MAD settings. However, in Meta-Judge settings, expanding the judge pool seems more beneficial than adding a bias-free agent, which are consistent with our reported findings. More results can be found in Table 16 and Table 17.
Judge Critic Round Pos.
4o-mini PINE 1 0.790
4o-mini PINE 2 0.792
4o-mini PINE 3 0.797
Table 16: MTbench results with Judge 4o-mini and Critic PINE.
Meta Judge Judgement Pool Choose Mode Conclude Mode
4o-mini 4o-mini, PINE 0.814 0.782
4o-mini R1-Distilled-Qwen32B, PINE 0.805 0.797
4o-mini Llama3.370B, R1-Distilled-Qwen32B, PINE 0.809 0.807
4o-mini DeepSeek-V3, Llama3.370B, R1-Distilled-Qwen32B, PINE 0.818 0.864
4o-mini 4o-mini, DeepSeek-V3, Llama3.370B, R1-Distilled-Qwen32B, PINE 0.814 0.824
Table 17: Results of Meta-Judge with PINE included in the Judgement pool.
Appendix D More information on Benchmarks
D.1 MT-Bench
@@
The first benchmark that we use in our work is MT-Bench Zheng et al. (2023), which originally contains responses from 30 LLMs to 80 questions. The questions are divided into eight categories: Writing, Roleplay, Reasoning, Math, Coding, Extraction, STEM, and Humanities, with ten questions in each category. The LLMs include GPT-4 Turbo, Claude Instant v1, Vicuna 13B v1.3, Alpaca 13B, and others. To construct our evaluation set, we treat each question paired with two distinct model responses as a single pairwise comparison task. For each question, we generate all possible unique response pairs between the selected LLMs, ensuring no repetition. From this set, we randomly sample 60 response pairs per category, each coupled with its corresponding question. This procedure yields a total of 480 comparison tasks across the eight categories. MT-bench is public available with License Apache 2.0.
D.2 CALM Benchmark
@@
Appendix E More information on PINE
@@
Appendix F Prompt Examples
F.1 Multi-Agent-Debate
@@
F.2 LLM-as-Meta Judge Prompts
@@ F.3 Bias Prompts
Appendix G Case Studies of Multi-Agent Debate
G.1 Original Debate
@@ G.3 With PINE
Appendix H Case Studies of LLM-as-Meta-Judge
H.1 Original Debate
@@ H.3 With PINE
Appendix I Implementation and Reproducibility
@@
To ensure the reproducibility of our experiments, we set fixed random seeds in our codes. All language models are evaluated with a temperature of $\text{temperature}=0.01$ to minimize randomness in their responses.